ABC Company, a manufacturing firm, has supplied the following information from its accounting


Records for the year 2020


Purchases of raw materials Br.76,000


Direct labor cost 52,500


Supplies used 5,300


Factory insurance 1,050


Commissions paid 7,500


Factory supervision 9,675


Advertising 2,400


Material handling 11,000


Work-in-process inventory, January 1, 2020 47,500


Work-in-process inventory, December 31, 2020 42,000


Materials inventory, January 1, 2020 10,400


Materials inventory, December 31, 2020 28,500


Finished goods inventory, January 1, 2020 20,055


Finished goods inventory, December 31, 2020 10,750


Sales as of December 31, 2020 250,000


Sales as of December 31, 2020


Required


1. Prepare a statement of cost of goods manufactured. (10 points)


2. Prepare a statement of cost of goods sold. (5 points)


3. Prepare income statement (5-points)

1
Expert's answer
2022-01-18T09:59:29-0500

Cost of Goods mannufactured

Direct Materials+ Direct Labour+ Manufacturing Overhead+ Opening WIP- Ending WIP

Direct Materials

Beginning Raw materials= 10,400

Purchases of R/M = 76000

Ending R/M =(28500)

=57,900

Direct Labour = 52500

Manufacturing cost( Material Handling cost)= 11,000

Opening WIP = 47,500

Closing WIP =(42,000)

=5,500

"\therefore COGM=57,900+52,500+11,000+5,500=126,900"

Cost of Goods Sold

= Opening stock of finished goods+Purchases-Closing stock of finished goods

Opening stock = 20,055

Purchases( COGM)=126,900

Closing Stock = (10750)

=136,205

Income statement

Sales =250,000

Cost of goods sold =(136,205)

Gross profit =113,795